Fermi National Laboratory


Accelerator Update for August 12, 2005 - August 15, 2005

Friday, August 12
The day shift began with Operations monitoring store 4328 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and the Meson MCenter and MTest experiments taking beam.

At 9:58 AM, a Main Injector (MI) RF station (MIRF3) tripped off for the first of seven times in three hours and forty-five minutes.

Operations put the Recycler into shot setup at 10:03 AM to stash antiprotons.

The stashing ended and Pbar resumed stacking at 11 AM.

MCenter and MTest resumed taking beam at 11:10 AM.

At 1:16 PM, Operation reported that MIRF3 was still causing problems. An expert investigated and made adjustments..

At 8:44 PM, Operations reported losing the NuMI beam permit due to beam problems for the first of eight times in two hours.

Operations terminated store 4328 at 10:04 PM.

Operations prepared D0 for an access that began at 10:26 PM.

Saturday, August 13
The midnight shift began with Operations monitoring the Recycler stash, with Pbar stacking, with MiniBooNE and NuMI taking beam, and with MCenter and MTest taking beam.

D0 came out of access at 12:48 AM.

At 1 AM, Operations reported that a Booster RF station (BRF9) tripped off and wouldn't reset. Operators bypassed the station.

Operations put the Tevatron (TeV) into shot setup at 1:51 AM.

At 3:33 AM, Operations reported that the Recycler RF structure disappeared after the second transfer of antiprotons. Operators couldn't make the third transfer.

Operations established combination store 4330 at 4:11 AM with an initial luminosity of 97.95E30. Pbar resumed stacking and MCenter and MTest resumed taking beam.

At 4:31 AM, Operations reported that the MI, NuMI, and Pbar all coincidentally tripped off due to different reasons: MI due to beam position, NuMI due to beam in the kicker gap, and Pbar due to an AP1 power supply. This happened three more times in nine minutes.

Operations reported at 4:43 AM that the lost of the Pbar beam permit was the cause of the other two trips. Operators investigated.

Operations put the Recycler into shot setup at 9:02 AM to stash antiprotons. The stash ended at 10:01 AM and Pbar resumed stacking.

At 12:17 PM, an RF expert reported that BRF9 had an open cable somewhere in the tunnel and would need an access to repair.

The Recycler began preparing to stash antiprotons at 2:34 PM. The stash ended and Pbar resumed stacking at 3:16 PM.

MCenter and MTest resumed taking beam at 3:30 PM.

At 7:14 PM, Operations reported that BRF12 tripped off on a ground fault indication. When it tripped, many other Booster RF stations sparked. Operators turned stations off, turned intensities down, turned stations back on and up slowly, and managed to get everything back to normal intensities within five minutes.

Experimenters at MTest said at 9:23 PM that they were through with beam for the night.

Sunday, August 14
The midnight shift began with Operations monitoring store 4330 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations terminated store 4330 at 8:07 AM.

Operations prepared D0 for an access at 8:17 AM. They came out of access at 11:58 AM.

Operations put the TeV into shot setup at 12:43 PM.

Operations established combination store 4332 at 2:54 PM with an initial luminosity of 127.86E30.

Pbar resumed stacking at 3:01 PM but at a poor rate. By five minutes later, stacking had begun to slowly improve.

MCenter and MTest resumed taking beam at 3:11 PM.

At 5:13 PM, Operations reported that the Pbar core cooling system was off. Operators investigated and then called in an expert. The expert reported at 8:10 PM that a 24-volt power supply was bad.

At 8:23 PM, a Pbar expert reported that an access was needed to fix a power supply. Operations put Pbar and the Recycler into shot setup to stash antiprotons and had the stack transferred by 9:09 PM.

Operations prepared Pbar for an access at 9:18 PM. The experts replaced a fuse, repaired a PLC problem, and worked on a magnet stand. They were out of access by 11:01 PM.

Stacking resumed at 11:33 PM.

Monday, August 15
The midnight shift began with Operations monitoring store 4332 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, MCenter, and MTest taking beam.

MTest experimenters went home for the night at 12:51 AM.

At 1:59 AM, BRF7 tripped off. Operators investigated and found the station with no power. Operators called in an expert who repaired a power supply and had the station back up by 3:07 AM.

Pbar resumed stacking at 3:12 AM and MCenter resumed taking beam a bit later.

Operations prepared the Recycler to stash antiprotons at 5:26 AM.

At 6:03 AM, Operations reported losing a transfer of antiprotons, and then another at 6:13 AM.

The Recycler stash ended and Pbar resumed stacking at 6:25 AM.

At 7:54 AM, Pbar's lithium lens tripped off. It initially reset, but then tripped off again. An operator called in an expert.

The Plans for Monday
The plan for today is to fix the lithium lens. The store will stay in the TeV until late this evening or even later depending on how long it takes to repair the lithium lens.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/15/2005   email Fermilab